# The Effect of Omitted Variables on

### v3

@@ 1 Introduction
In this paper, we study the relationship between these two kinds of breakdown points in coefficient stability analyses used to assess the importance of omitted variables in linear regression analysis. These analyses examine how coefficients change when additional regressors are added, and have a long history in empirical economics. Folk wisdom holds that if the additional covariates increase the R-squared substantially, but the coefficient nonetheless does not change much, then we can be more confident that any further omitted variable bias is small. For example, see the textbook discussion on pages 74–78 of Angrist and Pischke (2015). This idea was first formalized by Altonji et al. (2005a, b, 2008). It was later substantially extended by Oster (2019) (hereafter Oster), an extremely influential paper with about 4500 Google Scholar citations as of December 2024. Finkelstein et al. (2021, AER, page 2706) describe it as providing “the now-standard methodology” and “the standard approach” to adjusting for selection on unobservables. [^3] Other papers on regression sensitivity include Mauro (1990), Murphy and Topel (1990), Frank (2000), Imbens (2003), Altonji et al. (2005a), Clarke (2009), Bellows and Miguel (2009), Hosman et al. (2010), Gonzalez and Miguel (2015), Krauth (2016), Cinelli and Hazlett (2020), and Diegert et al. (2022) for linear models and Rosenbaum and Rubin (1983), Rosenbaum (1995, 2002), Robins et al. (2000), Altonji, Elder, and Taber (2005a, 2008), Masten and Poirier (2018, 2020), and Altonji et al. (2019) for nonlinear models. Moreover, from 2019–2021, 23 papers published in the top five economics journals formally compute and report the results from this method, while another 8 papers informally reference this method in support of their analysis. Based on the data collected by Blandhol et al. (2022), this is more than the number of papers that use 2SLS per year, among papers published in top five economics journals. We discuss in detail how Oster’s (2019) results are used in empirical economics in appendix LABEL:sec:inPractice.
@@
In section 2 we review Oster’s (2019) method. The main sensitivity parameter for this method, denoted by $\delta$, is commonly interpreted as the ratio of the magnitude of selection on unobservables to the magnitude of selection on observables. We then study breakdown points for this sensitivity parameter in section 3. Oster’s Proposition 2 characterizes the explain away breakdown point. Following her recommendation, empirical researchers now routinely report estimates of this explain away breakdown point as a measure of robustness to omitted variables. We show that this explain away breakdown point is very different from the sign change breakdown point. In fact, we show that any time the explain away breakdown point is large—suggesting that omitted variables may be unimportant—a much smaller value of her sensitivity parameter can actually change the sign of the parameter of interest. In particular, we prove that the sign change breakdown point is bounded above by $\delta=1$. This is a substantial concern since 1 is often viewed as the cutoff for a robust result. Consequently, if we maintain 1 as the cutoff, our result implies that no empirical results are robust to sign changes, using Oster’s method. Put differently: The true causal effect can have the opposite sign as the baseline estimand any time the magnitude of selection on unobservables is at least as large as the magnitude of selection on observables.
@@
Concretely, in this example the baseline estimate is on the vertical intercept at $\delta=0$. This estimate is positive. How robust is this estimate to the presence of omitted variables? In the left plot we see that we must go all the way to $\delta=-32$ before the value $\beta=0$ is consistent with the data and assumptions. This is the number that researchers commonly report as a measure of robustness. However, from the right plot, we see that we only need to go to $\delta=0.586$ to find a value of $\beta$ that is negative. This value $\delta=0.586$ is the sign change breakdown point. Oster (2019) does not discuss this breakdown point, its companion Stata package psacalc does not compute it, and none of the top five papers we survey in appendix LABEL:sec:inPractice report it. Yet, as we see here, the sign change breakdown point is substantially smaller than $|-32|$, the explain away breakdown point. Indeed, it is below $\delta=1$, the commonly used cutoff for a robust result. Consequently, in this empirical example, focusing on the sign change breakdown point overturns the authors’ conclusion of robustness: Only a small amount of selection on unobservables relative to observables is necessary to flip the sign of the baseline estimate.
Importantly, all of the coefficient values plotted in figure 1, even the seemingly extreme values, are equally consistent with the data and assumptions. Under the assumptions of Oster’s Proposition 2, which is used to obtain these plots, there is no justification for ignoring any of these values. The only way to eliminate values is by making additional assumptions. One reasonable assumption, which we consider in section 3.2, is to restrict the magnitude of the omitted variable bias. With this additional assumption, we can justifiably eliminate some of the most extreme values. This assumption does not overturn our main conclusions, however. For example, in figure 1 the coefficient value consistent with $\delta=0.586$ is not extreme at all—it is the same order of magnitude as the baseline estimate. So in this example, the sign change breakdown point is still very different from the explain away breakdown point, even with a substantial restriction on the magnitude of the omitted variable bias. Moreover, magnitude restrictions require researchers to choose a tuning parameter—just how much should they restrict the magnitude? If the magnitude is restricted too much, then this assumption is essentially equivalent to assuming a priori that there is no omitted variable bias, in which case a sensitivity analysis would be unnecessary. But as soon as reasonably sized magnitudes are allowed, the concerns we discuss in sections 3 and 4 arise.
@@
To address the issues described in sections 3 and 4, we propose several simple modifications to the methods in Oster (2019). First, we recommend gathering all values of the coefficient consistent with a range of sensitivity parameters, and presenting this set. The left plot in figure 2 gives an example: Here we gather all $\beta$’s consistent with $|\delta|$ less than some fixed $\bar{\delta}$. We then plot that set as a function of $\bar{\delta}$. The graph of these values of $\beta$ is discontinuous in $\bar{\delta}$ at $\bar{\delta}=0.586$, which explains the difference in the sign change breakdown and explain away breakdown points. As we saw in figure 1, nearly all values of $\beta$, including arbitrarily large negative values, are consistent with an upper bound of $\bar{\delta}=1.05$, but a small range around 0 is not part of this set until $\bar{\delta}$ equals 32. Researchers who want to focus only on the largest and smallest possible values of $\beta$ can instead present the convex hull, which is shown in the right plot of figure 2.
@@
In section 6 we illustrate our results using data from Satyanath et al. (2017, Journal of Political Economy), who study the effect of social capital on the rise of the Nazi party. There we show just how different the sign change and explain away breakdown points can be in practice. We then show how magnitude restrictions can sometimes help retain findings of robustness. We also illustrate the extreme sensitivity of bias adjusted estimates to perturbations of the sensitivity parameter. In appendix LABEL:sec:extraEmpirical we study three additional empirical applications in detail, all published in the American Economic Review.
To see that these four applications are not outliers, in section 7 we conduct two meta-analyses covering 58 different empirical applications published in the top five economics journals plus the AEJ: Applied Economics. We show that the distinction between explain away and sign change breakdown points leads to substantively different conclusions about robustness to omitted variables in the majority of empirical applications. In particular, suppose researchers were to continue the current standard practice of reporting explain away breakdown points, but wanted to interpret them as sign change breakdown points. We show that in more than 50% of applications, this interpretation is only valid if researchers already know the sign of coefficient of interest. But in that case, it would not be necessary to do a sensitivity analysis, since the result is known a priori.
@@ 2.1 Regressions of Interest
Suppose $\beta_{\text{long}}$ is the parameter of interest. In particular, we can write the long regression as
@@ 2.2 Comparing Selection on Observables and on Unobservables
$\delta$ is the ratio of these two regression coefficients. $\delta$ is not known from the data since it depends on $W_{2}$, which is not observed, and $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$, which are also unknown. Instead, we will study what can be said about $\beta_{\text{long}}$ under assumptions on $\delta$. Several papers propose alternative sensitivity parameters, including Cinelli and Hazlett (2020) and Diegert et al. (2022). In this paper we instead use the same sensitivity parameter $\delta$ as Oster (2019), since our focus here is to study the implications of using different kinds of breakdown points. We give more discussion of $\delta$’s interpretation in section 5 below.
@@
$\delta$ is defined to calibrate the magnitude of selection on unobservables by using the observed covariates $W_{1}$. Typically we will have additional covariates $W_{0}$ which we want to include in the analysis, but not use for calibration. Denote these baseline controls by $W_{0}$. To incorporate these covariates, replace $(X,W_{1},W_{2})$ with $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$ throughout the entire analysis. See section 3.3 of Diegert et al. (2022) for a justification of this approach and further discussion. For simplicity, we omit $W_{0}$ throughout our theoretical analysis.
@@ 2.3 Identification of $\beta_{\text{long}}$
A4 requires that all of the observed covariates are uncorrelated with the omitted variable. Appendix A.1 in Oster (2019) briefly discusses one approach to relaxing A4, which we discuss in section 5. Also see Appendix A of Diegert et al. (2022) for additional details. Diegert et al. (2022) also develop an alternative approach to sensitivity analysis that does not rely on A4. In this paper we maintain A4 throughout because our main focus is to study different notions of breakdown. For simplicity, from here on we assume $W_{2}$ is a scalar. All of our results can be generalized to vector $W_{2}$ using the arguments in Appendix B of Diegert et al. (2022).
@@ 2.4 Oster’s Breakdown Analysis
Theorem 1 is very similar to Oster’s Proposition 3, with only minor differences. As Oster remarked, this result shows that, even though knowledge of $\delta$ does not point identify $\beta_{\text{long}}$, knowledge of $\beta_{\text{long}}$ does point identify $\delta$. Figure 1 illustrates why: Fixing $\delta$ is equivalent to placing a vertical line on the plot. This vertical line may intersect the displayed curve up to three times, since $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$ can have up to three elements. Fixing the value $\beta_{\text{long}}$ to be equal to some known $\beta_{\text{hypo}}$, however, is equivalent to placing a horizontal line on the plot. This line can only intersect the displayed curve once. The absolute value of the horizontal coordinate at this intersection is precisely $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$.
@@ 3.1 The Sign Change Breakdown Point for $\delta$
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ that is required for the data to be consistent with $\beta_{\text{long}}=0$?
@@
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $\delta$ such that $\beta_{\text{long}}$ has a different sign from $\beta_{\text{med}}$?
@@
Here we assume $\beta_{\text{short}}\neq\beta_{\text{med}}$ for simplicity only. Theorem 2 shows that, in Oster’s sensitivity analysis, the sign change breakdown point can never be larger than one. In practice, researchers typically consider the value $\delta=1$ to be the cutoff for determining the robustness of a result. That is, they interpret values of $\delta$ greater than 1 as robust and values less than 1 as sensitive. Based on this criterion, Theorem 2 therefore shows that the sign of the parameter of interest is never robust (assuming we require $\delta$ strictly larger than 1 to conclude that the result is robust). In contrast, the explain away breakdown point $\delta^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ is often substantially larger than one. We illustrate this difference empirically in section 6.
@@
    $\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}.$
    $\gamma_{2,\text{long}}$, as one of the long regression coefficients, will also be large because of near multi-collinearity. Consequently, this measure of selection on unobservables will be close to zero since regression coefficients scale down when the covariate is scaled up.
@@
Thus in the first step we see that the OVB will be large while in the third step we see that $\delta$ will be close to one. This is precisely what happens along the asymptote at $\delta=1$. This followed from (a) analyzing the behavior of near multi-collinear OLS coefficients and (b) exploring the specific structure of $\delta$. Thus the asymptote arises because $(X,W_{1},W_{2})$ are allowed to be arbitrarily mutli-collinear. Researchers may be interested in explicitly ruling out this collinearity. This is exactly what the alternative methods of Cinelli and Hazlett (2020) and Diegert et al. (2022) do. Both papers include sensitivity parameters that explicitly bound the joint explanatory power of the observed and unobserved covariates in the first stage.

### v4

@@ 1 Introduction
In this paper, we study the relationship between these two kinds of breakdown points in coefficient stability analyses used to assess the importance of omitted variables in linear regression analysis. These analyses examine how coefficients change when additional regressors are added, and have a long history in empirical economics. Folk wisdom holds that if the additional covariates increase the R-squared substantially, but the coefficient nonetheless does not change much, then we can be more confident that any further omitted variable bias is small. For example, see the textbook discussion on pages 74–78 of Angrist and Pischke (2015). This idea was first formalized by Altonji et al. (2005a, b, 2008). It was later substantially extended by Oster (2019) (hereafter Oster), an extremely influential paper with about 4900 Google Scholar citations as of June 2025. Finkelstein et al. (2021, AER, page 2706) describe it as providing “the now-standard methodology” and “the standard approach” to adjusting for selection on unobservables. [^3] Other papers on regression sensitivity include Mauro (1990), Murphy and Topel (1990), Frank (2000), Imbens (2003), Altonji et al. (2005a), Clarke (2009), Bellows and Miguel (2009), Hosman et al. (2010), Gonzalez and Miguel (2015), Krauth (2016), Cinelli and Hazlett (2020), Diegert et al. (2025a, b) for linear models and Rosenbaum and Rubin (1983), Rosenbaum (1995, 2002), Robins et al. (2000), Altonji, Elder, and Taber (2005a, 2008), Masten and Poirier (2018, 2020), and Altonji et al. (2019) for nonlinear models. Moreover, from 2019–2021, 23 papers published in the top five economics journals formally compute and report the results from this method, while another 8 papers informally reference this method in support of their analysis. Based on the data collected by Blandhol et al. (2022), this is more than the number of papers that use 2SLS per year, among papers published in top five economics journals. We discuss in detail how Oster’s (2019) results are used in empirical economics in Appendix LABEL:sec:inPractice.
@@
In Section 2 we review Oster’s (2019) method. The main sensitivity parameter for this method, denoted by $\delta$, is commonly interpreted as the ratio of the magnitude of selection on unobservables to the magnitude of selection on observables. We then study breakdown points for this sensitivity parameter in Section 3. Oster’s Proposition 2 characterizes the explain away breakdown point. Following her recommendation, empirical researchers now routinely report estimates of this explain away breakdown point as a measure of robustness to omitted variables. We show that this explain away breakdown point is very different from the sign change breakdown point. In fact, we show that any time the explain away breakdown point is large—suggesting that omitted variables may be unimportant—a much smaller value of her sensitivity parameter can actually change the sign of the parameter of interest. In particular, we prove that the sign change breakdown point is bounded above by $1$. This is a substantial concern since 1 is often viewed as the cutoff for a robust result. Consequently, if we maintain 1 as the cutoff, our result implies that no empirical results are robust to sign changes, using Oster’s method. Put differently: The true causal effect can have the opposite sign as the baseline estimand any time the magnitude of selection on unobservables is at least as large as the magnitude of selection on observables.
@@
Concretely, the baseline estimate is obtained at $\delta=0$. In this empirical example, the baseline estimate is positive. How robust is this estimate to the presence of omitted variables? In the left plot we see that we must go all the way to $\delta=-32$ before the value $\beta=0$ is consistent with the data and assumptions. This is the number that researchers commonly report as a measure of robustness. However, from the right plot, we see that we only need to go to $\delta=0.586$ to find a value of $\beta$ that is negative. This value $|\delta|=0.586$ is the sign change breakdown point. Oster (2019) does not discuss this breakdown point, its companion Stata package psacalc does not compute it, and none of the papers we survey in Appendix LABEL:sec:inPractice report it. Yet, as we see here, the sign change breakdown point is substantially smaller than $|-32|$, the explain away breakdown point. Indeed, its magnitude is below $1$, the commonly used cutoff for a robust result. Consequently, in this empirical example, focusing on the sign change breakdown point overturns the authors’ conclusion of robustness: Only a small amount of selection on unobservables relative to observables is necessary to flip the sign of the baseline estimate.
Importantly, all of the coefficient values plotted in Figure 1, even the seemingly extreme values, are equally consistent with the data and assumptions. Under the assumptions of Oster’s Proposition 2, which is used to obtain these plots, there is no justification for ignoring any of these values. The only way to eliminate values is by making additional assumptions. One reasonable assumption, which we consider in Section 3.2, is to restrict the magnitude of the omitted variable bias. With this additional assumption, we can justifiably eliminate some of the most extreme values. This assumption does not overturn our main conclusions, however. For example, in Figure 1 the coefficient value consistent with $\delta=0.586$ is not extreme at all—it is the same order of magnitude as the baseline estimate. So in this example, the sign change breakdown point is still very different from the explain away breakdown point, even with a substantial restriction on the magnitude of the omitted variable bias. Moreover, magnitude restrictions require researchers to choose a tuning parameter—just how much should they restrict the magnitude? If the magnitude is restricted too much, then this assumption is essentially equivalent to assuming a priori that there is no omitted variable bias, in which case a sensitivity analysis would be unnecessary. But as soon as reasonably sized magnitudes are allowed, the concerns we discuss in sections 3 and 4 can arise.
@@
To address the issues described in sections 3 and 4, we propose several simple modifications to the methods in Oster (2019). First, we recommend gathering all values of the coefficient consistent with a range of sensitivity parameters, and presenting this set. The left plot in Figure 2 gives an example: Here we gather all $\beta$’s consistent with $|\delta|$ less than some fixed $\bar{\delta}$. We then plot that set as a function of $\bar{\delta}$. The graph of these values of $\beta$ is discontinuous in $\bar{\delta}$ at $\bar{\delta}=0.586$, which explains the difference in the sign change breakdown and explain away breakdown points. As we saw in Figure 1, nearly all values of $\beta$, including arbitrarily large negative values, are consistent with an upper bound of $\bar{\delta}=1.05$, but a small range around 0 is excluded from the set, and the value 0 does not become part of this set until $\bar{\delta}$ equals 32. Researchers who want to focus only on the largest and smallest possible values of $\beta$ can instead present the convex hull, which is shown in the right plot of Figure 2.
@@
In Section 6 we illustrate our results using data from Satyanath et al. (2017, Journal of Political Economy), who study the effect of social capital on the rise of the Nazi party. There we show just how different the sign change and explain away breakdown points can be in practice. We then show how magnitude restrictions can sometimes help retain findings of robustness. We also illustrate the extreme sensitivity of bias adjusted estimates to perturbations of the sensitivity parameter. In Appendix LABEL:sec:extraEmpirical_applications we study three additional empirical applications in detail, all published in the American Economic Review.
To see that these four applications are not outliers, in Section 7 we conduct two meta-analyses covering 58 different empirical applications published in the top five economics journals plus the AEJ: Applied Economics. We show that the distinction between explain away and sign change breakdown points leads to substantively different conclusions about robustness to omitted variables in the majority of empirical applications. In particular, suppose researchers were to continue the current standard practice of reporting explain away breakdown points, but wanted to interpret them as sign change breakdown points. We show that in more than 50% of applications, this interpretation is only valid if researchers already know the sign of the coefficient of interest before performing the sensitivity analysis. But in that case, it would not be necessary to do a sensitivity analysis, since the result is known a priori.
@@ 2.1 Regressions of Interest
It is well known that linear regression coefficients can be given causal interpretations under various identifying assumptions (e.g., Angrist and Pischke 2009). In this paper, we take it as given that $\beta_{\text{long}}$ is the parameter of interest. This focus on $\beta_{\text{long}}$ is often because we have in fact made assumptions that make it a causal parameter, but our technical results do not rely on those motivating assumptions. So, taking interest in $\beta_{\text{long}}$ as given, write the long regression as
@@ 2.2 Comparing Selection on Observables and on Unobservables
$\delta$ is the ratio of these two regression coefficients. $\delta$ is not known from the data since it depends on $W_{2}$, which is not observed, and $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$, which are also unknown. Instead, we will study what can be said about $\beta_{\text{long}}$ under assumptions on $\delta$. Several papers propose alternative sensitivity parameters, including Cinelli and Hazlett (2020) and Diegert et al. (2025a). In this paper we instead use the same sensitivity parameter $\delta$ as Oster (2019), since our focus here is to study the implications of using different kinds of breakdown points. We give more discussion of $\delta$’s interpretation in Section 5 below.
@@
$\delta$ is defined to calibrate the magnitude of selection on unobservables by using the observed covariates $W_{1}$. Typically we will have additional covariates $W_{0}$ which we want to include in the analysis, but not use for calibration. Denote these baseline controls by $W_{0}$. To incorporate these covariates, replace $(X,W_{1},W_{2})$ with $(X^{\perp W_{0}},W_{1}^{\perp W_{0}},W_{2}^{\perp W_{0}})$, the variables after residualizing with respect to $W_{0}$, throughout the entire analysis. See Section 3.4 of Diegert et al. (2025a) for a justification of this approach and further discussion. For simplicity, we omit $W_{0}$ throughout our theoretical analysis.
@@ 2.3 Identification of $\beta_{\text{long}}$
A4 requires that all of the observed covariates are uncorrelated with the omitted variable. Appendix A.1 in Oster (2019) briefly discusses one approach to relaxing A4, which we discuss in Section 5. Also see Appendix C of Diegert et al. (2025b) for additional details. Diegert et al. (2025a) develop an alternative approach to sensitivity analysis that does not rely on A4. In this paper we maintain A4 throughout because our main focus is to study different notions of breakdown. For simplicity, from here on we assume $W_{2}$ is a scalar. All of our results can be generalized to vector $W_{2}$ by working with the scalar index $\gamma_{2,\text{long}}^{\prime}W_{2}$ everywhere (as in the proof of Theorem 2 in Diegert et al. 2025a, for example).
@@ 2.4 Oster’s Breakdown Analysis
The proof of Theorem 1 and all other formal results can be found in Appendix LABEL:sec:proofs. Theorem 1 is very similar to Oster’s Proposition 3, with only minor differences. As Oster remarked, this result shows that, even though knowledge of $\delta$ does not point identify $\beta_{\text{long}}$, knowledge of $\beta_{\text{long}}$ does point identify $\delta$. Figure 1 illustrates why: Fixing $\delta$ is equivalent to placing a vertical line on the plot. This vertical line may intersect the displayed curve up to three times, since $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})$ can have up to three elements. Fixing the value $\beta_{\text{long}}$ to be equal to some known $\beta_{\text{hypo}}$, however, is equivalent to placing a horizontal line on the plot. This line can only intersect the displayed curve once. The absolute value of the horizontal coordinate at this intersection is precisely $\delta^{\text{bp},=}(\beta_{\text{hypo}},R_{\text{long}}^{2})$.
@@ 3.1 The Sign Change Breakdown Point for $\delta$
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $|\delta|$ that is required for the data to be consistent with $\beta_{\text{long}}=0$?
@@
  • Suppose we knew the exact value of $R_{\text{long}}^{2}$. What is the smallest value of $|\delta|$ such that $\beta_{\text{long}}$ has a different sign from $\beta_{\text{med}}$?
@@
Here we assume $\beta_{\text{short}}\neq\beta_{\text{med}}$ for simplicity only. Theorem 2 shows that, in Oster’s sensitivity analysis, the sign change breakdown point can never be larger than one. In practice, researchers typically consider the value $|\delta|=1$ to be the cutoff for determining the robustness of a result. That is, they interpret values of $|\delta|$ greater than 1 as robust and values less than 1 as sensitive. Based on this criterion, Theorem 2 therefore shows that the sign of the parameter of interest is never robust (assuming we require $|\delta|$ strictly larger than 1 to conclude that the result is robust). In contrast, the explain away breakdown point $\delta^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ is often substantially larger than one. We illustrate this difference empirically in Section 6.
@@
    $\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2})}.$ (6)
    $\gamma_{2,\text{long}}$, as one of the long regression coefficients, will also be large because of near multi-collinearity. Consequently, this measure of selection on unobservables will be close to zero since regression coefficients scale down when the covariate is scaled up. More formally, note that $\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2})$ and $\operatorname*{var}(\gamma_{2,\text{long}}W_{2})$ both diverge to infinity as we approach perfect multi-collinearity since $|\gamma_{2,\text{long}}|\rightarrow\infty$ in this scenario. However their ratio (the regression coefficient in (6)) converges to zero since its numerator is linear in $\gamma_{2,\text{long}}$ while its denominator is quadratic in $\gamma_{2,\text{long}}$.
@@
Thus in the first step we see that the OVB will be large while in the third step we see that $\delta$ will be close to one. This is precisely what happens along the asymptote at $\delta=1$. This followed from (a) analyzing the behavior of near multi-collinear OLS coefficients and (b) exploring the specific structure of $\delta$. Thus the asymptote arises because $(X,W_{1},W_{2})$ are allowed to be arbitrarily mutli-collinear. Researchers may be interested in explicitly ruling out this collinearity. This is exactly what the alternative methods of Cinelli and Hazlett (2020) and Diegert et al. (2025a) do. Both papers include sensitivity parameters that explicitly bound the joint explanatory power of the observed and unobserved covariates in the first stage. The sensitivity analyses proposed by both papers then lead to sign change and explain away breakdown points that are equal.